OpenStudy (anonymous):

express using a positive exponent y^-4= would the answer be 1/y^4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Great thank you!!

OpenStudy (anonymous):

36c^2+1-12c factoring completely would that be (1-6c)^2

OpenStudy (anonymous):

right again

OpenStudy (anonymous):

Awesome thanks so much again!!

OpenStudy (anonymous):

\[\sqrt[5]{-\left(\begin{matrix}1 \\ 243\end{matrix}\right)}\] Simplify

OpenStudy (anonymous):

Is that a matrix/vector or are you trying to show division?

OpenStudy (anonymous):

Its a fraction